(Real options and capital budgeting​) You have come up with a great idea for a​ Tex-Mex-Thai fusion restaurant. After doing a financial analysis of this​ venture, you estimate that the initial outlay will be $6 million. You also estimate that there is a 50 percent chance that this new restaurant will be well received and will produce annual cash flows of $760,000 per year forever​ (a perpetuity), while there is a 50 percent chance of it producing a cash flow of only $200,000 per year forever​ (a perpetuity) if it​ isn't received well.

a. What is the NPV of the restaurant if the required rate of return you use to discount the project cash flows is 10 ​percent?

b. What are the real options that this analysis may be​ ignoring?

c. Explain why the project may be worthwhile even though you have just estimated that its NPV is​ negative?
Answer #1

Solution (a)

Given

Two scenarios with equal probability:

Scenario 1 (Well received): Future cash flows of $760,000 in perpetuity

Scenario 2 (not well received): Future cash flows of $200,000 in perpetuity

probability of each scenario = 0.50

initial outlay = $6 million

discount rate for future cash flows = 10% (0.10 in decimal)

Formula

Present Value (PV) of perpetuity = (Annual cash flow) / (discount rate in decimal)

NPV = (PV of perpetuity) - (initial outlay)

Project NPV = (Probability of Scenario 1 * Scenario 1 NPV) +  (Probability of Scenario 2 * Scenario 2 NPV)

Scenario 1 NPV

PV of $760,000 in perpetuity = (Annual cash flow) / (discount rate in decimal) = $ (760,000) / (0.10) = $7,600,000 = $7.6 million

Hence, NPV = (PV of perpetuity) - (initial outlay) = $ (7.6 - 6) million = $1.6 million

Scenario 2 NPV

PV of $200,000 in perpetuity =  (Annual cash flow) / (discount rate in decimal) = $ (200,000) / (0.10) = $2,000,000 = $2 million

Hence, NPV = (PV of perpetuity) - (initial outlay) = $ (2 - 6) million = $(-)4 million

Therefore, NPV of entire project = (Probability of Scenario 1 * Scenario 1 NPV) +  (Probability of Scenario 2 * Scenario 2 NPV) = (0.50 * 1.6) + ( 0.50 * - 4) = $ (0.80 - 2) = $(-)1.2 million


Answer: NPV of restaurant is $(-)1.2 million.

Solution (b)

The real options being ignored are the situations in which:

(i) Actions could be taken to change the value proposition and improve the future cash flows even in case the restaurant is not well received.

(ii) Option to expand or contract or abandon can be excercised based on changing environment or market dynamics.

Solution (c)


Even though the estimated NPV is negative, the project may be pursued in case:

(I) The discount rate of 10% used is higher compared to the risk involved or the opportunity cost of capital.

(ii) The restaurant could be financed at a rate lower than the discount rate assumed to calculate NPV.

(iii) There is no alternative option of similar risk where a positive NPV can be obtained.

(iv) Other factors such as payback period is favourable/manageable.

(v) There are real options likely to be available in future to boost cash flows or lower risk and hence discount rate.